a baker has modeled the monthly operating costs for making wedding cakes by the function y=0.5x^2-12+150 where y is the total cost in dollars and x is the number of cakes prepared. what is the vertex and axis of symmetry?
Answer by stanbon(75887)

a baker has modeled the monthly operating costs for making wedding cakes by the function y=0.5x^2-12x+150 where y is the total cost in dollars and x is the number of cakes prepared. what is the vertex and axis of symmetry?
----
Vertex occurs when x = -b/(2a) = 12/(2*0.5) = 12
f(12) = (1/2)(144) -12*12 +150 = -72+150 = 78
----
Vertex at (12,78)
Axis of symmetry: x = 12
=============================
Cheers,
Stan H.
=============